# A Diffusion Model to Shrink Proteins While Maintaining Their Function

### v1

@@ 5 Using SCISOR to generate and shrink protein sequences
Note that SCISOR is not simply sampling deletions by how natural $\mathrm{prev}(X)$ looks. Rather it also uses knowledge of $M$ to plan for future mutations. Different values of $M$ allow the model to change which deletions it will allow at each step (Fig. 3b).
Given enough data and compute, SCISOR should learn the correct distribution, $q_{\theta}(X_{0}=\tilde{X}\mid X,M)$. However, in principle, there is a distribution shift between sampling $q_{\theta}(X_{0}=\tilde{X}\mid X,M)$ when $X$ is a realistic protein versus when $X$ is a sequence with noisy insertions in our learning process, which may impact the statistical efficiency of the learning process. In practice, App. C and our results in the following sections show that the SCISOR de-noiser learns meaningful evolutionary signals.
@@ 8 Preserving key in-silico indicators of function when shrinking
Since the diffusion models in Sec. 6 cannot suggest deletions, we compare as baselines to shrinking with ProGen2 and Raygun. Raygun requires $1$ model evaluation to make $M$ deletions, while SCISOR requires $M$. For ProGen2, ideally we would sample substrings of $X$ with length LML-M proportional to ProGen2’s likelihood; however, that approach would require $\binom{L}{M}$ model evaluations, which is prohibitively expensive. We instead consider a strong but tractable baseline: we predict the effect of all $L$ single deletions independently (in $L$ model evaluations) and then sample sets of deletions with probabilities based on these effects. Further details are in appendix˜B and section˜E.3.
In Fig. 6 we see SCISOR consistently suggests shrunken proteins that are more likely to be foldable and preserve key indicators of function (predicted structural topology, presence of active or binding sites) than Raygun and ProGen2 baselines. Although ProGen2 achieves higher functional site preservation when shrinking by 50%, the lower pLDDT scores suggest that its samples at this level of shrinkage are not likely to be functional. In App. E.2, we show that SCISOR also achieves best performance for greedy sampling of shrunk sequences, reflecting situations in which a practitioner is not interested in generating diverse samples. Lastly, in appendix˜D we perform an in-depth case study of shrinking the GTP sensor RalA, showing that shrinking with SCISOR best preserves the predicted structure of the binding site with GTP.
@@ B.3 Training SCISOR
For the results in section˜6, we train SCISOR models on the March 2020 release of UniRef50, using the same train-test split as EvoDiff [Alamdari2023-nj] from https://zenodo.org/records/6564798. Our models were trained about one week each on one NVIDIA A100 GPU with an effective batch size of 256 and learning rate of 0.0001.
For the results in sections˜8 and 7, we train SCISOR models on the latest release of UniRef90. Here, we use an effective batch size of 512 and learning rate of 0.00005. The SCISOR S and M models were trained for about one week each on two NVIDIA A100 GPUs. The SCISOR L model was trained for about four days on four NVIDIA H100 GPUs.
@@ B.6.2 Model evaluation
Appendix C SCISOR deletions match those in nature for R4SNK4
We collected 5000 sequences locally aligned to R4SNK4 using BLAST [Camacho2009-og] and counted how frequently each position of the R4SNK4 sequence landed in a region that wasn’t aligned (“Not aligned”) or aligned with a gap token (“Gap”). These are two distinct ways to measure deletions made by nature. Next, we shrunk R4SNK4 using SCISOR at various values of $M$, taking 512 samples for each $M$.
In Fig. 7a, we plot, for each position of R4SNK4, how frequently we observe deletions as suggested by SCISOR versus as seen in nature. We see qualitatively that at $M=50$, SCISOR prefers making deletions in regions where deletions are observed in nature, and at $M=100$ SCISOR prefers deleting the second half of the protein, the region which was less frequently included in alignments. This separate behavior at different $M$ is shown quantitatively when we measure the Spearman correlations between SCISOR deletion frequencies and deletions seen in the alignments in Fig. 7b. Almost all comparisons are statistically significant – a random model never achieves a Spearman correlation above 0.015.
Figure 7: SCISOR suggests deletions in R4SNK4 seen in nature. (a) We compare two ways to measure deletions seen in an alignment versus deletions at two values of $M$ from samples of SCISOR. (b) We quantitatively measure the Spearman correlations between the two types of deletions in our alignment to deletions seen in SCISOR samples across $M$.
@@ E.2 Deterministic Shrinking
In section˜8, we report results for sampling shrunk sequences from SCISOR and baseline models. Those evaluations are most relevant to practical settings, where practitioners may wish to generate a diverse set of promising candidate designs to be evaluated in the lab. Here, we instead consider deterministic shrinking, where the most likely deletions are performed. For Raygun, we mimic this setting by using setting the noise parameter to 0. In fig.˜8, we see that SCISOR consistently achieves excellent performance in this setting as well. Although ProGen2 can achieve higher performance in some cases, this method is more computationally expensive than SCISOR. Moreover, our metrics may be rewarding the “myopic” nature of our ProGen2 baseline, which likely favors deletions that preserve functional regions in the original protein that are not necessarily functional in the shrunken protein.
Figure 8: Deterministically shrinking proteins with SCISOR yields high-quality shrunk samples. All details are as in fig.˜6, except that we take the most likely deletions predicted from each model. For Raygun, we set $\textrm{noise}=0$.
@@ E.3 Computationally Expensive ProGen2 Baseline
In section˜8, we reported a baseline for ProGen2 that required $L$ forward passes of the model, already exceeding the $M$ forward passes used by SCISOR. That baseline assumes that the effects of each deletion are independent, which may be unrealistic. In table˜10, we present results for a more expensive ProGen2 baseline, requiring $\mathcal{O}(L\cdot M)$ forward passes, where deletions are performed one at a time, and the effects of subsequent deletions recalculated. We see that switching to this more expensive ProGen2 baseline does not significantly improve performance.
% Deleted ProGen2 - $\mathcal{O}(L)$ ProGen2 - $\mathcal{O}(LM)$ SCISOR - $\mathcal{O}(M)$
1 71.6 71.7 74.0
3 68.2 67.9 74.2
5 63.5 62.8 74.1
10 55.4 53.0 73.1
20 40.7 36.4 71.2
30 32.3 31.2 63.7
40 31.3 31.9 55.1
50 30.9 33.5 44.7
Table 10: Average pLDDT structural confidence scores from Omegafold for various amounts of shrinkage.
@@ F.1 Proof of correctness for Algorithm 1
using the same logic as above. Therefore, $\mathrm{Var}_{q}\mathrm{ali}(X_{0},X)=o\left(\binom{M+L}{L}^{2}\right)=o(E_{q}\mathrm{ali}(X_{0},X)^{2})$. This completes the proof.
@@ F.6 Useful lemma
$p(X_{t}\mid X_{0},M_{t})=$ $\sum_{\text{indices}\ i_{1},\dots,i_{L}}\frac{\mathbbm{1}(X_{0}=X_{t}^{(i_{1})}\cdots X_{1}^{(i_{l})})}{\binom{L+M_{t}}{L}}\times\prod_{b\in X_{t}\setminus X_{0}}\pi(b)$
$=$ $\binom{L+M_{t}}{L}^{-1}\mathrm{ali}(X_{0},X_{t})\prod_{b\in X_{t}\setminus X_{0}}\pi(b).$

### v2

@@ 5 Using SCISOR to generate and shrink protein sequences
Note that SCISOR is not simply sampling deletions by how natural $\mathrm{prev}(X)$ looks. Rather it also uses knowledge of $M$ to plan for future mutations. Different values of $M$ allow the model to change which deletions it will allow at each step (Fig. 3b). Given enough data and compute, SCISOR should learn the correct distribution $q_{\theta}(X_{0}=\tilde{X}\mid X,M)$. In principle however, there is a distribution shift between sampling $q_{\theta}(X_{0}=\tilde{X}\mid X,M)$ when $X$ is a realistic protein and our learning process when $X$ is a sequence with noisy insertions; this may impact the statistical efficiency of the learning process. In practice, App. C and our results in the following sections show that the SCISOR de-noiser learns meaningful evolutionary signals. We also confirm this using a toy example in App. H.
@@ 6 SCISOR is a competitive generative model for proteins
In E.5, we ablate several components of our framework, particularly highlighting the importance of our Rao-Blackwellized training objective.
@@ 8 Preserving key in-silico indicators of function when shrinking
Since the diffusion models in Sec. 6 cannot suggest deletions, we compare as baselines to shrinking with ProGen2 and Raygun. Raygun requires $1$ model evaluation to make $M$ deletions, while SCISOR requires $M$. For ProGen2, ideally we would sample substrings of $X$ with length LML-M proportional to ProGen2’s likelihood; however, that approach would require $\binom{L}{M}$ model evaluations, which is prohibitively expensive. We instead consider a strong but tractable baseline: we predict the effect of all $L$ single deletions independently (in $L$ model evaluations) and then sample sets of deletions with probabilities based on these effects. Further details are in App. B and App. E.3.
In Fig. 6 we see SCISOR consistently suggests shrunken proteins that are more likely to be foldable and preserve key indicators of function (predicted structural topology, presence of active or binding sites) than Raygun and ProGen2 baselines. Although ProGen2 achieves higher functional site preservation when shrinking by 50%, the lower pLDDT scores suggest that its samples at this level of shrinkage are not likely to be functional. We stratify these results by Omegafold pLDDT, ESM2 pseudo-likelihood, and cellular localization in App. E.4 and App. E.6. In App. E.2, we show that SCISOR also achieves best performance for greedy sampling of shrunk sequences, reflecting situations in which a practitioner is not interested in generating diverse samples. Lastly, in App. D we perform an in-depth case study of shrinking the GTP sensor RalA, showing that shrinking with SCISOR best preserves the predicted structure of the binding site with GTP.
@@ B.3 Training SCISOR
For the results in Sec. 6, we train SCISOR models on the March 2020 release of UniRef50, using the same train-test split as EvoDiff [Alamdari2023-nj] from https://zenodo.org/records/6564798. Our models were trained about one week each on one NVIDIA A100 GPU with an effective batch size of 256 and learning rate of 0.0001.
For the results in Sec. 8 and Sec. 7, we train SCISOR models on the latest release of UniRef90. Here, we use an effective batch size of 512 and learning rate of 0.00005. The SCISOR S and M models were trained for about one week each on two NVIDIA A100 GPUs. The SCISOR L model was trained for about four days on four NVIDIA H100 GPUs.
@@ B.6.2 Model evaluation
Appendix C SCISOR deletions match those in nature
We collected sequences locally aligned to three proteins using the phmmer web server [Finn2011-qg] (open penalty=0.01, Extend=0.2, and all other settings default) and counted how frequently each position of each sequence landed in a region that wasn’t aligned (“Not aligned”) or aligned with a gap token (“Gap”). These are two distinct ways to measure deletions made by nature. Next, we shrunk each sequence using SCISOR at various values of $M$, taking 512 samples for each $M$.
We ran our analysis on a catalyst R4SNK4_9PSEU, a membrane transporter TPIS_HUMAN, and a transcription factor FOXA1_HUMAN.
In Fig. 7, we plot, for each position of each protein, how frequently we observe deletions as suggested by SCISOR versus as seen in nature. We see strong correlation between the samples from SCISOR and those from nature. Almost all comparisons are statistically significant – a random model never achieves a Spearman correlation above 0.015, while we achieve a p-value of $10^{-10}$ for each protein.
Interestingly, for R4SNK4_9PSEU we observe separate behavior at different $M$: we see qualitatively that at $M=50$, SCISOR prefers making deletions in regions where gaps are observed in nature, while at $M=100$ SCISOR prefers deleting large chunks of the protein, the region which was less frequently included in alignments.
(a) R4SNK4_9PSEU
(c) TPIS_HUMAN
(e) FOXA1_HUMAN
Figure 7: SCISOR suggests deletions seen in nature. (a, c, e) We compare two ways to measure deletions seen in an alignment versus deletions at two values of $M$ from samples of SCISOR. (b, d, f) We quantitatively measure the Spearman correlations between the two types of deletions in our alignment to deletions seen in SCISOR samples across $M$.
@@ E.2 Deterministic Shrinking
In Sec. 8, we report results for sampling shrunk sequences from SCISOR and baseline models. Those evaluations are most relevant to practical settings, where practitioners may wish to generate a diverse set of promising candidate designs to be evaluated in the lab. Here, we instead consider deterministic shrinking, where the most likely deletions are performed. For Raygun, we mimic this setting by using setting the noise parameter to 0. In Fig. 8, we see that SCISOR consistently achieves excellent performance in this setting as well. Although ProGen2 can achieve higher performance in some cases, this method is more computationally expensive than SCISOR. Moreover, our metrics may be rewarding the “myopic” nature of our ProGen2 baseline, which likely favors deletions that preserve functional regions in the original protein that are not necessarily functional in the shrunken protein.
Figure 8: Deterministically shrinking proteins with SCISOR yields high-quality shrunk samples. All details are as in Fig. 6, except that we take the most likely deletions predicted from each model. For Raygun, we set $\textrm{noise}=0$.
@@ E.3 Computationally Expensive ProGen2 Baseline
In Sec. 8, we reported a baseline for ProGen2 that required $L$ forward passes of the model, already exceeding the $M$ forward passes used by SCISOR. That baseline assumes that the effects of each deletion are independent, which may be unrealistic. In Fig. 9, we present results for a more expensive ProGen2 baseline, requiring $\mathcal{O}(L\cdot M)$ forward passes, where deletions are performed one at a time, and the effects of subsequent deletions recalculated. We see that switching to this more expensive ProGen2 baseline does not improve performance, and SCISOR remains the best-performing method.
Figure 9: SCISOR outperforms expensive ProGen2 baseline. All details are as in Fig. 6, except that we include a more expensive ProGen2 baseline that requires $\mathcal{O}(L\cdot M)$ forward passes compared to SCISOR’s $\mathcal{O}(M)$.
E.4 Shrinking Results Stratified by Wild Type pLDDT and ESM2 Pseudo-Likelihood
In Fig. 10, we show how SCISOR’s shrunk designs compare to the baselines, stratified by the Omegafold pLDDT quartile of the wild type. SCISOR consistently produces high-quality designs across all settings, confirming its robustness.
Figure 10: SCISOR consistently predicts high-quality shrunk designs for wild type proteins with various pLDDT scores. All details are as in Fig. 6, and we facet results by quartile of wild type pLDDT score, as predicted by Omegafold.
In Fig. 11, we show how SCISOR’s shrunk designs compare to the baselines, stratified by the ESM2 pseudo-likelihood quartile of the wild type, similar to the analysis in Gordon2024-nh. The pseudo-likelihood is calculated by masking one index at a time, computing the likelihood of that element, and then aggregating the scores across the full sequence. SCISOR consistently produces high-quality designs across all settings, confirming its robustness.
Figure 11: SCISOR consistently predicts high-quality shrunk designs for wild type proteins with various pLDDT scores. All details are as in Fig. 6, and we facet results by quartile of wild type pLDDT score, as predicted by Omegafold.
E.5 Ablations of modelling choices
We perform an ablation study of training SCISOR S on Uniref50, ablating two aspects of our model. In Fig. 12a, we see that our structured forward process which weights amino acids by their prevalence results in a slight speedup of training. In contrast, in Fig. 12b, we see that the Rao-Blackwellization described in Sec. 4.2 is crucial in enabling the model to learn efficiently.
Figure 12: Our structured forward process and Rao-Blackwellization of the loss enable efficient training of SCISOR. We train SCISOR S on Uniref50 sequences, ablating components of our framework. While (a) switching from a uniform distribution over amino acids to one that weights amino acids by their prevalence results in a small improvement, (b) the Rao-Blackwellized loss is crucial to efficient training of SCISOR. The random baseline is based on the entropy of the amino acid distribution in Uniref50. We plot an exponential moving average of the perplexity on a withheld validation set.
E.6 Features enriched in proteins that are hard to shrink
We investigated the differences between proteins in the analysis in Fig. 6, hypothesizing that membrane proteins may be more challenging to shrink. In Fig. 13 we find no enrichment in membrane proteins among the sequences whose structure dropped the most in quality after shrinking. This is potentially because they often contain large regions outside of the membrane, or because intra-membrane regions are not particularly sensitive to deletions. A preliminary analysis did reveal a slight enrichment of extra-cellular proteins however.
Figure 13: We do not have evidence membrane proteins are harder to shrink. We plot the results of Fig. 6 for 200 proteins. We plot the pLDDT of the full-length protein against that shrunk 20% by SCISOR. We colour membrane and extra-cellular proteins.
@@ F.1 Proof of correctness for Algorithm 1
using the same logic as above. Therefore, $\mathrm{Var}_{q}\mathrm{ali}(X_{0},X)=o\left(\binom{M+L}{L}^{2}\right)=o(E_{q}\mathrm{ali}(X_{0},X)^{2})$. This completes the proof. ∎
@@ F.6 Useful lemma
$\begin{aligned} p(X_{t}\mid X_{0},M_{t})=&\sum_{\text{indices}\ i_{1},\dots,i_{L}}\frac{\mathbbm{1}(X_{0}=X_{t}^{(i_{1})}\cdots X_{1}^{(i_{l})})}{\binom{L+M_{t}}{L}}\times\prod_{b\in X_{t}\setminus X_{0}}\pi(b)\\ =&\binom{L+M_{t}}{L}^{-1}\mathrm{ali}(X_{0},X_{t})\prod_{b\in X_{t}\setminus X_{0}}\pi(b).\end{aligned}\qed$
@@ Appendix G Alignments algorithm
Appendix H Experiment with Toy Synthetic Dataset
There is a distribution shift between the sequences SCISOR sees during training - composed of unnatural sequences with potentially many random insertions - and the inputs we wish to use when shrinking natural proteins. To show that this distribution shift is not a fundamental issue for SCISOR, we perform an experiment with a toy synthetic dataset designed to capture this challenge.
Specifically, our toy synthetic data has an alphabet of A, B, C, but assumes that the only functional sequences are those that alternate A and B, such as “ABABABABA”. The length of functional sequences is sampled uniformly at random between 1 and 20.
During training, due to the random insertions, SCISOR is exposed mostly to sequences that are not valid, such as “ACBBABCCBAB”. Nevertheless, when we apply SCISOR to shrink a valid sequence, it should learn to only return valid subsequences. Indeed, in Fig. 14, we see that this is the case: within half an hour of training, SCISOR samples >>99% functional subsequences from functional sequences. All training details are the same as for our protein experiments, except we use a smaller batch size of 64.
(a) SCISOR trains on out-of-distribution data
(b) SCISOR learns to shrink valid sequences
Figure 14: Despite being trained on unnatural sequences, SCISOR is able to shrink natural sequences while preserving validity. (a) We train SCISOR S on a toy synthetic dataset where natural sequences are alternating As and Bs, such that the random insertions noising process results in primarily invalid sequences. (b) We find that despite being trained to shrink invalid sequences, SCISOR quickly learns to shrink valid sequences as well in less than 30 minutes of training. Here, we plot the percent of alternating sequences of length $L=20$ which SCISOR shrinks to another valid sequence after $M=10$ deletions.